A heavy-duty truck produces a 50Hz sound wave. If the speed of sound in air is 336 m/s, what is the wavelength of this sound wave.


1
Expert's answer
2020-11-30T14:52:47-0500

We can find the wavelength of the sound wave from the wave speed formula:


"v=f\lambda,""\lambda=\dfrac{v}{f}=\dfrac{336\ \dfrac{m}{s}}{50\ Hz}=6.72\ m."

Answer:

"\lambda=6.72\ m."